1. Problem 3.34 (a) (c).

2. Problem 3.39

3. Probelm 3.41 (b) (e)

4. Problem 3.55

5. Consider a system with the impulse response

h(n) = (\frac{1}{2}) cos(ω_{o}n)u(n).

(a) Find H(z).

(b) Plot the poles and zeros of H(z).

(c) Write the difference equation that characterizes the system.

(d) Is the system stable? Is it causal?